[Hallc_running] [New Logentry] RC: Daily update

mdiefent at jlab.org mdiefent at jlab.org
Thu Dec 13 09:55:02 EST 2018

Logentry Text:
<p><strong><span style="font-family: Helvetica; font-size: 12px;">Update on swing shift of December 12 and owl shift of December 13</span></strong><br style="font-family: Helvetica; font-size: 12px;"><br style="font-family: Helvetica; font-size: 12px;"><strong><span style="font-family: Helvetica; font-size: 12px;">CEBAF status: </span></strong></p><ul><li><span style="font-family: Helvetica; font-size: 12px;">Received 7.6 hours of beam with currents up to 70 uA during the swing shift (https://logbooks.jlab.org/entry/3638213). </span></li><li><span style="font-family: Helvetica; font-size: 12px;">Received 5.6 hours of beam with a current of 13 uA during the owl shift (https://logbooks.jlab.org/entry/3638317). There have been occasional IIC3H04 trips. </span></li></ul><div style="font-family: Helvetica; font-size: 12px;"><strong>Hall C status: </strong><br><ul><li>Three-pass beam with an energy of 4.933 GeV is delivered to Hall C. </li><li>Hall C is setup for t
 he Kaon LT experiment (E12-09-011). The experiments aims for the first precision measurements of L/T separated kaon electroproduction cross sections above the resonance region. Details on the Kaon LT experiment can be found on: https://redmine.jlab.org/projects/kltexp/wiki</li><li>On Monday, December 17, we will change during a control access the SHMS angle to angle of 6.00 degree. </li></ul></div><div style="font-family: Helvetica; font-size: 12px;"><strong>Data taking: </strong><div><ul><li>The detailed run plan is available on: https://redmine.jlab.org/attachments/download/590/runplan_4p9gev.pdf</li><li>We have done the harp scan and taken the data for the calibration runs. </li><li>The beam energy measurement is scheduled for tomorrow at the beginning of the day shift. </li><li>We have taken the data for the calibration runs. </li><li>We have started data taking in the high epsilon setting. The run conditions allows us to take data for bea
 m currents of 13uA, 20uA, 26uA. </li></!
 ul></div><div> </div><div><strong>Hall C Daily RC meeting</strong>: We will meet today at 4:15 p.m. (EDT) in the conference room on the second floor of the counting house. Remote participation will be possible via BlueJeans (ID: 861439920). You can join via:  <div><ul><li>browser: https://bluejeans.com/861439920</li><li>phone: (888) 240-2560 (US toll free); list of international phone numbers: http://bluejeans.com/number</li></ul></div></div><div>We will review the status of the experiment, discuss the analysis of the calibration runs, and coordinate the work for the next days. </div></div>


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3638365
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20181213/098e87bf/attachment-0001.html>

More information about the Hallc_running mailing list